# A Persistence-Driven Edit Distance for Graphs with Abstract Weights

### v3

@@ 6.2.1 Notation
Consider $x\in V_{T}$ and $y\in V_{T^{\prime}}$. Along with keeping the notation defined in Section 6.1, define $T_{x}:=sub_{T}(x)$ and $T_{y}:=sub_{T^{\prime}}(y)$, $N_{x}:=\dim(T_{x})=\#E_{T}$ and $N_{y}:=\dim(T_{y})=\#E_{T^{\prime}}$. Moreover, given $v\in V_{T_{x}}$, the sequence $v_{0}=v<v_{1}<\ldots<r_{T}$ indicates the points in the segment $\zeta_{v}$. Thus $v_{i}$ will be a vertex $v_{i}>v$. The same with $w\in V_{T_{y}}$
@@ 6.2.2 Relaxing the Optimization Problem
We would like to find $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ minimizing Equation (2) for $T_{x}$ and $T_{y}$, but this is a difficult task. In fact, as evident in the construction of $\alpha(M^{*})$, a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ has the role of pairing segments: if $(v,w)\in M^{*}$, then the segments $\zeta_{v}^{\Omega(v)}$ and $\zeta_{w}^{\Omega(w)}$ are shrunk one on the other by $\alpha(M^{*})$. However, the points $\Omega(v)$ and $\Omega(w)$ depend on the whole set $M^{*}$, and not simply on the couple $(v,w)$. Modelling such global dependence gives rise to non-linear relationships between coupled points, and so leading to a non linear cost function, in terms of points interactions, to be minimized. For this reason we “weaken” the last term in Equation (2), allowing also mappings different from $\alpha(M^{*})$ to be built from $M^{*}$. In other words we minimize over $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ the following equation:
@@
  • $\beta(M^{*})\in M_{\mathcal{G}}(\widetilde{T}_{x},\widetilde{T}_{y})$ (using the notation established in 7, replacing $T$ and $T^{\prime}$ with $T_{x}$ and $T_{y}$ respectively);
  • the set of vertices coupled by $\beta(M^{*})$ is exactly $M^{*}$: $M^{*}=\beta(M^{*})\bigcap V_{T_{x}}\times V_{T_{y}}$.
Since, by construction $M^{*}=\alpha(M^{*})\bigcap V_{T}\times V_{T^{\prime}}$ and by 7, $\alpha(M^{*})\in M_{2}(\widetilde{T}_{x},\widetilde{T}_{y})$, minimizing Equation (2) or Equation (3) gives the same results.
@@ 6.2.3 Setup and Variables
The mapping $\beta(M^{*})$ is built according to the variables $\delta^{v,w}_{i,j}$ with value $1$: we write a constrained optimization problem such that having $\delta^{v,w}_{i,j}=1$ means pairing the segments $\zeta_{v}^{v_{i+1}}$ (that is, the segment which starts with $(v,v_{1})$ and ends with $(v_{i},v_{i+1})$) and $\zeta_{w}^{w_{j+1}}$, and shrinking one in the other in the induced mapping.
In order to pair and shrink the segments $\zeta_{v}^{v_{i+1}}=\{v=v_{0},v_{1},\ldots,v_{i+1}\}$ and $\zeta_{w}^{w_{j+1}}$ we need to collect some edits in the set $\beta(M^{*})$ adding the following edits:
@@ 6.2.4 Constraints
Clearly, the set of $\delta^{v,w}_{i,j}$ with value $1$ does not always produce $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$: for instance segments could be paired multiple times. To avoid such issues, we build a set of constraints for the variable $\delta$.
@@
  • the couples $(v,w)$ such that $\delta^{v,w}_{i,j}=1$ define a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$;
  • the edits induced by all $\delta^{v,w}_{i,j}=1$ give a mapping $\beta(M^{*})$ in $M_{``G^{\prime\prime}}(\widetilde{T}_{x},\widetilde{T}_{y})$. With $\widetilde{T}_{x},\widetilde{T}_{y}$ being obtained from $T_{x}$ and $T_{y}$ as in 7.
@@ 6.2.5 Objective Function
As before, consider $v\in E_{T_{x}}$ and interpret $\delta^{v,w}_{i,j}=1$ as coupling the segments $\zeta_{v}^{v_{i+1}}$ and $\zeta_{w}^{w_{j+1}}$; then $v$ is coupled with some $w\in E_{T_{y}}$ if $C(v):=\sum_{i,w,j}\delta^{v,w}_{i,j}=1$ and is ghosted if $G(v):=\sum_{\{i,v^{\prime}\mid v\in\zeta_{v^{\prime}}^{v^{\prime}_{i}}\}}\sum_{w,j}\delta^{v^{\prime},w}_{i,j}=1$. The vertex $v$ is instead deleted if $D(v):=1-C(v)-G(v)=1$. We introduce also the following quantities, which correspond to the cost of shrinking $\zeta_{v}^{v_{i+1}}$ on $\zeta_{w}^{w_{j+1}}$:
@@
The function which computes the cost given by coupled points is therefore:
@@
Finally, one must take into account the values of $d_{E}(T_{v},T_{w})$, whenever $v$ and $w$ are coupled; this information is contained in $(W_{xy})_{v,w}$:
@@
Proof  The contribution of coupled points is $F^{C}(\delta)$ and the contribution of deleted points is $F^{D}(\delta)-F^{-}(\delta)$.
The cost of $\beta(M^{*})$ is: $F^{\beta}(\delta):=F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)$. Lastly, $F^{S}(\delta)$ takes into account the value of $d_{E}(T_{v},T_{w})$, if $v$ and $w$ are coupled. By 3, combined with 8, the solution of the following optimization problem:
@@
6.3 Computing the Edit Distance for Trees: Bottom-Up Algorithm
@@
Given $x\in V_{T}$, define $\operatorname*{len}(x)$ to be the number of vertices in $\zeta_{x}$ and $\operatorname*{len}(T)=\max_{v\in V_{T}}\operatorname*{len}(v)$. Therefore, $\operatorname*{lvl}(x)=\operatorname*{len}(T)-\operatorname*{len}(x)$. Lastly, $\operatorname*{lvl}_{T}(n)=\{v\in V_{T}\mid\operatorname*{lvl}(v)=n\}$
The key property is that: $\operatorname*{lvl}(x)>\operatorname*{lvl}(v)$ for any $v\in sub(x)$. Thus, if $W_{xy}$ is known for any $x\in\operatorname*{lvl}_{T}(n)$ and $y\in\operatorname*{lvl}_{T^{\prime}}(m)$, then for any $v$, $w$ in $V_{T}$, $V_{T^{\prime}}$ such that $\operatorname*{lvl}(v)<n$ and $\operatorname*{lvl}(w)<m$, $W_{vw}$ is known as well. With this notation we can write down Algorithm 1.
@@ 6.4 Computational Cost
Proof  In a full binary tree structure, at each level $l$ we have $2^{l}$ vertices. Let $L=\operatorname*{len}(T)$ and $L^{\prime}=len(T^{\prime})$. We have that, for any vertex $v\in V_{T}$ at level $l$, the cardinality of the path from $v$ to any of the leaves in $sub_{T}(v)$ is ${L-l}$ and the number of leaves in $sub_{T}(v)$ is $2^{L-l}$.
@@
We end the manuscript with a remark about the computational cost of $d_{E}$ as described in 10. We found this result surprising to some extent. If we consider the latest implementation of the classical edit distance between unlabelled and unordered trees obtained with BLP (Hong et al., 2017), then the complexity of the two metrics is very similar: the classical edit distance can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot M)$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints - $O(N+M)$ if we count also the constraints restricting the integer variables to $\{0,1\}$, as the authors of Hong et al. (2017) do. Thus, the introduction of the ghosting and splitting edits, with all the complexities they carry, just increases the cost by two $\log$ factors in the number of variables. Moreover, the condition used in Zhang (1996) to make the classical edit distance tractable (that is, item (2) in Section 3.1. of Zhang (1996)) is compatible with (M1)-(M4) and thus a constrained version of $d_{E}$ is well defined. We believe that it could be computed with a polynomial time algorithm. We leave this investigation to future works.
@@ 7 Conclusions
We develop a novel framework to work with graphs that arise as topological summaries in the context of topological data analysis. These kind of summaries are increasingly frequent and require some particular care to be properly handled by a metric. We define an edit distance employing some novel edit operations, proving that such definition induces a metric up to a certain equivalence relationship between graphs.
For situations in which the graphs possess the structure of a tree then, by exploiting the ordering properties of trees, we produce a binary linear programming algorithm to compute such metric. The modifications that we make w.r.t. the classical edit distance between unlabeled trees make it only marginally more expensive, namely by some log factors. For this reason we argue that polynomial time approximations should be available also for $d_{E}$. Still, the computational procedure we describe here is used in separate applied works, as mentioned in the introduction. These additional modifications are essential for the stability properties assessed in Pegoraro (2024a), justifying the higher computational cost w.r.t. other edit distances for merge trees, which, instead, are unstable.
@@
  • Reeb Graphs can be investigated theoretically and algorithmically, perhaps following the decomposition presented in Stefanou (2020);
  • likewise, computational procedures and theoretic investigations should be considered for particular finite sub-posets for $\mathbb{R}^{n}$;
  • as motivated by 11 we believe that polynomial time approximations can be obtained for computing $d_{E}$ between dendrograms;
@@ Acknowledgments
This work was carried out as part of my PhD Thesis, under the supervision of Professor Piercesare Secchi.
@@ A Algorithmic Example
We consider the following couple of weighted trees. Let $(T,w_{T})$ be the tree given by: $V_{T}=\{a,b,c,d,r_{T}\}$, $E_{T}=\{(a,d),(b,d),(d,r_{T}),(c,r_{T})\}$ and $w_{T}(a)=w_{T}(b)=w_{T}(d)=1$, $w_{T}(c)=5$; the tree $(T^{\prime},w_{T^{\prime}})$ instead, is defined by: $V_{T^{\prime}}=\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$, $E_{T^{\prime}}=\{(a^{\prime},d^{\prime}),(b^{\prime},d^{\prime}),(d^{\prime},r_{T^{\prime}}),(c^{\prime},r_{T^{\prime}})\}$ and $w_{T^{\prime}}(a)=1$, $w_{T^{\prime}}(b)=w_{T^{\prime}}(c)=2$ and $w_{T^{\prime}}(d)=3$. We employ the notation $w_{T}$ for the weight function instead of $\varphi_{T}$ to indicate the common situation of a tree with positive real weights.
@@
This step is trivial since we only have couples between leaves, like $(a,a^{\prime})$, which have trivial subtrees and thus $d_{E}(sub_{T}(a),sub_{T^{\prime}}(a^{\prime}))=0$.
@@
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 1$ are $\{a,b,c,d\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 1$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime}\}$. Thus the couples $(x,y)$ which are considered are: $(d,d^{\prime})$, $(d,a^{\prime})$, $(d,b^{\prime})$, $(d,c^{\prime})$ and $(a,d^{\prime})$, $(b,d^{\prime})$, $(c,d^{\prime})$. The couples between leaves, like $(a,a^{\prime})$ have already been considered.
Couple: $(d,d^{\prime})$
@@
Couple: $(d,a^{\prime})$
Obviously: $d_{E}(sub_{T}(d),sub_{T^{\prime}}(a^{\prime}))=\mid\mid sub_{T}(d)\mid\mid$. All the couples featuring a leaf and an internal vertex (that is, a vertex which is not a leaf), such as $(d,b^{\prime})$, $(a,d^{\prime})$ etc. behave similarly.
@@
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 2$ are $\{a,b,c,d,r_{T}\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 2$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$. Thus the couples $(x,y)$ which are considered are $(d,r_{T^{\prime}})$, $(r_{T},d^{\prime})$, $(r_{T},r_{T^{\prime}})$ and then the trivial ones: $(r_{T},a^{\prime})$, $(r_{T},b^{\prime})$, $(r_{T},c^{\prime})$ and $(a,r_{T^{\prime}})$, $(b,r_{T^{\prime}})$, $(c,r_{T^{\prime}})$. Some couples have already been considered and thus are not repeated.
Couple: $(d,r_{T^{\prime}})$
@@
Couple: $(r_{T},d^{\prime})$
This and the other couples are left to the reader.

### v4

@@ 6.2.1 Notation
Consider $x\in V_{T}$ and $y\in V_{T^{\prime}}$. Along with keeping the notation defined in Section 6.1, define $T_{x}:=sub_{T}(x)$ and $T_{y}:=sub_{T^{\prime}}(y)$, $N_{x}:=\dim(T_{x})=\#E_{T}$ and $N_{y}:=\dim(T_{y})=\#E_{T^{\prime}}$. Moreover, given $v\in V_{T_{x}}$, the sequence $v_{0}=v<v_{1}<\ldots<r_{T}$ indicates the points in the path $\zeta_{v}$. Thus $v_{i}$ will be a vertex $v_{i}>v$. The same with $w\in V_{T_{y}}$
@@ 6.2.2 Relaxing the Optimization Problem
We would like to find $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ minimizing Equation (3) for $T_{x}$ and $T_{y}$, but this is a difficult task. In fact, as evident in the construction of $\alpha(M^{*})$, a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ has the role of pairing paths: if $(v,w)\in M^{*}$, then the paths $\zeta_{v}^{\Omega(v)}$ and $\zeta_{w}^{\Omega(w)}$ are shrunk one on the other by $\alpha(M^{*})$. However, the points $\Omega(v)$ and $\Omega(w)$ depend on the whole set $M^{*}$, and not simply on the pair $(v,w)$. Modeling such global dependence gives rise to non-linear relations between paired points, and so leading to a non linear cost function, in terms of points interactions, to be minimized. For this reason we “weaken” the last term in Equation (3), allowing also mappings different from $\alpha(M^{*})$ to be built from $M^{*}$. In other words we minimize over $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ the following equation:
@@
  • $\beta(M^{*})\in M_{\mathcal{G}}(\widetilde{T}_{x},\widetilde{T}_{y})$ (using the notation established in 6, replacing $T$ and $T^{\prime}$ with $T_{x}$ and $T_{y}$ respectively);
  • the set of vertices paired by $\beta(M^{*})$ is exactly $M^{*}$: $M^{*}=\beta(M^{*})\bigcap V_{T_{x}}\times V_{T_{y}}$.
Since, by construction $M^{*}=\alpha(M^{*})\bigcap V_{T}\times V_{T^{\prime}}$ and by 6, $\alpha(M^{*})\in M_{2}(\widetilde{T}_{x},\widetilde{T}_{y})$, minimizing Equation (3) or Equation (4) gives the same results.
@@ 6.2.3 Setup and Variables
The mapping $\beta(M^{*})$ is built according to the variables $\delta^{v,w}_{i,j}$ with value $1$: we write a constrained optimization problem such that having $\delta^{v,w}_{i,j}=1$ means pairing the paths $\zeta_{v}^{v_{i+1}}$ (that is, the path which starts with $(v,v_{1})$ and ends with $(v_{i},v_{i+1})$) and $\zeta_{w}^{w_{j+1}}$, and shrinking one in the other in the induced mapping.
In order to pair and shrink the paths $\zeta_{v}^{v_{i+1}}=\{v=v_{0},v_{1},\ldots,v_{i+1}\}$ and $\zeta_{w}^{w_{j+1}}$ we need to collect some edits in the set $\beta(M^{*})$ adding the following edits:
@@ 6.2.4 Constraints
Clearly, the set of $\delta^{v,w}_{i,j}$ with value $1$ does not always produce $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$: for instance paths could be paired multiple times. To avoid such issues, we build a set of constraints for the variable $\delta$.
@@
  • the pairs $(v,w)$ such that $\delta^{v,w}_{i,j}=1$ define a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$;
  • the edits induced by all $\delta^{v,w}_{i,j}=1$ give a mapping $\beta(M^{*})$ in $M_{``G^{\prime\prime}}(\widetilde{T}_{x},\widetilde{T}_{y})$. With $\widetilde{T}_{x},\widetilde{T}_{y}$ being obtained from $T_{x}$ and $T_{y}$ as in 6.
@@ 6.2.5 Objective Function
As before, consider $v\in E_{T_{x}}$ and interpret $\delta^{v,w}_{i,j}=1$ as pairing the paths $\zeta_{v}^{v_{i+1}}$ and $\zeta_{w}^{w_{j+1}}$; then $v$ is paired with some $w\in E_{T_{y}}$ if $C(v):=\sum_{i,w,j}\delta^{v,w}_{i,j}=1$ and is ghosted if $G(v):=\sum_{\{i,v^{\prime}\mid v\in\zeta_{v^{\prime}}^{v^{\prime}_{i}}\}}\sum_{w,j}\delta^{v^{\prime},w}_{i,j}=1$. The vertex $v$ is instead deleted if $D(v):=1-C(v)-G(v)=1$. We introduce also the following quantities, which correspond to the cost of shrinking $\zeta_{v}^{v_{i+1}}$ on $\zeta_{w}^{w_{j+1}}$:
@@
The function which computes the cost given by paired points is therefore:
@@
Finally, one must take into account the values of $d_{E}(T_{v},T_{w})$, whenever $v$ and $w$ are paired; this information is contained in $(W_{xy})_{v,w}$:
@@
Proof  The contribution of paired points is $F^{C}(\delta)$ and the contribution of deleted points is $F^{D}(\delta)-F^{-}(\delta)$.
The cost of $\beta(M^{*})$ is: $F^{\beta}(\delta):=F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)$. Lastly, $F^{S}(\delta)$ takes into account the value of $d_{E}(T_{v},T_{w})$, if $v$ and $w$ are paired. By 3, combined with 7, the solution of the following optimization problem:
@@
6.3 Bottom-Up Algorithm
@@
Given $x\in V_{T}$, define $\operatorname*{len}(x)$ to be the number of vertices in $\zeta_{x}$ and $\operatorname*{len}(T)=\max_{v\in V_{T}}\operatorname*{len}(v)$. Then, we set $\operatorname*{lvl}(x)=\operatorname*{len}(T)-\operatorname*{len}(x)$.
The key property is that: $\operatorname*{lvl}(x)>\operatorname*{lvl}(v)$ for any $v\in sub(x)$. Thus, if $W_{xy}$ is known for any $x\in\operatorname*{lvl}_{T}^{-1}(n)$ and $y\in\operatorname*{lvl}_{T^{\prime}}^{-1}(m)$, then for any $v$, $w$ in $V_{T}$, $V_{T^{\prime}}$ such that $\operatorname*{lvl}(v)<n$ and $\operatorname*{lvl}(w)<m$, $W_{vw}$ is known as well. With this notation we can write down Algorithm 1.
@@ 6.4 Computational Cost
Proof  In a full binary tree structure, at each level $l$ we have $2^{l}$ vertices. Let $L=\operatorname*{len}(T)$ and $L^{\prime}=\operatorname*{len}(T^{\prime})$. We have that, for any vertex $v\in\operatorname*{lvl}(l)^{-1}$, the cardinality of the path from $v$ to any of the leaves in $sub_{T}(v)$ is ${L-l}$ and the number of leaves in $sub_{T}(v)$ is $2^{L-l}$.
@@
We end the manuscript with a remark about the computational cost of $d_{E}$ as described in 9. We find this result surprising to some extent. If we consider a recent implementation of the classical edit distance between unlabeled and unordered trees obtained with BLP (Hong et al., 2017), then the complexity of the two metrics is very similar: the classical edit distance can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot M)$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints - $O(N+M)$ if we count also the constraints restricting the integer variables to $\{0,1\}$, as the authors of Hong et al. (2017) do. Thus, the introduction of the ghosting and splitting edits, with all the complexities they carry, just increases the cost by two $\log$ factors in the number of variables. Moreover, the condition used in Zhang (1996) to make the classical edit distance tractable (that is, item (2) in Section 3.1. of Zhang (1996)) is compatible with (M1)-(M4) and thus a constrained version of $d_{E}$ is well defined. We believe that it could be computed with a polynomial time algorithm. We leave this investigation to future works.
@@ 7 Conclusions
We develop a novel framework to work with tree that arise as topological summaries in the context of topological data analysis, possibly enriched with some abstract weights as a result of different pipelines. These kind of summaries are increasingly frequent and require some particular care to be properly handled by a metric. We define an edit distance employing some novel edit operations, proving that such definition induces a metric up to a certain equivalence relation between weighted trees.
By exploiting the ordering properties of trees, we produce a binary linear programming algorithm to compute such metric. The modifications that we make w.r.t. the classical edit distance between unlabeled trees make it only marginally more expensive, namely by some log factors. For this reason we argue that polynomial time approximations should be available also for $d_{E}$. Still, the computational procedure we describe here is used in separate applied works, as mentioned in the introduction. These additional modifications are essential for the stability properties assessed in Pegoraro and Secchi (2024); Pegoraro (2024b), justifying the higher computational cost w.r.t. other edit distances for merge trees, which, instead, are unstable.
@@
  • we would like to extend this edit distance outside tree-shaped graphs, encompassing Reeb graphs (Biasotti et al., 2008; De Silva et al., 2016) and their combinatorial approximation, namely mapper graphs (Singh et al., 2007) are studied (Carriere et al., 2018; Carriere and Oudot, 2018), perhaps following the decomposition presented in Stefanou (2020). We also mention that works on GEDs like (Ambauen et al., 2003; Lerouge et al., 2016) could be of great help in developing such generalization;
  • as motivated by 8 we believe that polynomial time approximations can be obtained for computing $d_{E}$ between dendrograms;
@@ Acknowledgments
This work was carried out as part of my PhD Thesis, under the supervision of Professor Piercesare Secchi. I also acknowledge the support of the Wallenberg AI, Autonomous Systems and Software Program (WASP), and of the SciLifeLab and Wallenberg National Program for Data-Driven Life Science (DDLS), which fund the project: Topological Data Analysis of Functional Genome to Find Covariation Signatures.
@@ A Algorithmic Example
We consider the following pair of weighted trees. Let $(T,w_{T})$ be the tree given by: $V_{T}=\{a,b,c,d,r_{T}\}$, $E_{T}=\{(a,d),(b,d),(d,r_{T}),(c,r_{T})\}$ and $w_{T}(a)=w_{T}(b)=w_{T}(d)=1$, $w_{T}(c)=5$; the tree $(T^{\prime},w_{T^{\prime}})$ instead, is defined by: $V_{T^{\prime}}=\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$, $E_{T^{\prime}}=\{(a^{\prime},d^{\prime}),(b^{\prime},d^{\prime}),(d^{\prime},r_{T^{\prime}}),(c^{\prime},r_{T^{\prime}})\}$ and $w_{T^{\prime}}(a)=1$, $w_{T^{\prime}}(b)=w_{T^{\prime}}(c)=2$ and $w_{T^{\prime}}(d)=3$. We employ the notation $w_{T}$ for the weight function instead of $\varphi_{T}$ to indicate the common situation of a tree with positive real weights.
@@
This step is trivial since we only have pairs between leaves, like $(a,a^{\prime})$, which have trivial subtrees and thus $d_{E}(sub_{T}(a),sub_{T^{\prime}}(a^{\prime}))=0$.
@@
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 1$ are $\{a,b,c,d\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 1$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime}\}$. Thus the pairs $(x,y)$ which are considered are: $(d,d^{\prime})$, $(d,a^{\prime})$, $(d,b^{\prime})$, $(d,c^{\prime})$ and $(a,d^{\prime})$, $(b,d^{\prime})$, $(c,d^{\prime})$. The pairs between leaves, like $(a,a^{\prime})$ have already been considered.
Pair: $(d,d^{\prime})$
@@
Pair: $(d,a^{\prime})$
Obviously: $d_{E}(sub_{T}(d),sub_{T^{\prime}}(a^{\prime}))=\mid\mid sub_{T}(d)\mid\mid$. All the pairs featuring a leaf and an internal vertex (that is, a vertex which is not a leaf), such as $(d,b^{\prime})$, $(a,d^{\prime})$ etc. behave similarly.
@@
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 2$ are $\{a,b,c,d,r_{T}\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 2$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$. Thus the pairs $(x,y)$ which are considered are $(d,r_{T^{\prime}})$, $(r_{T},d^{\prime})$, $(r_{T},r_{T^{\prime}})$ and then the trivial ones: $(r_{T},a^{\prime})$, $(r_{T},b^{\prime})$, $(r_{T},c^{\prime})$ and $(a,r_{T^{\prime}})$, $(b,r_{T^{\prime}})$, $(c,r_{T^{\prime}})$. Some pairs have already been considered and thus are not repeated.
Pair: $(d,r_{T^{\prime}})$
@@
Pair: $(r_{T},d^{\prime})$
This and the other pairs are left to the reader.